OpenStudy (anonymous):

PLEASE HELP ME!!! The function below shows the number of car owners f(t), in thousands, in a city in different years t: f(t) = 1.1t2 - 2.5t + 1.5 The average rate of change of f(t) from t = 3 to t = 5 is ______ thousand owners per year.

OpenStudy (godmod360):

9.9 - 7.5 + 1.5 =3.9

OpenStudy (anonymous):

So are you sure about that? =) @godmod360

OpenStudy (amorfide):

considering he only substituted in t=3 he has no idea what he is doing

OpenStudy (anonymous):

Okay so how would u propose doing it, @amorfide =)

OpenStudy (amorfide):

the question mentions rate of change, therefore we will differentiate f(t) then you would substitute t=5, t=3 and subtract the answers

OpenStudy (anonymous):

1.1(3)^2 -2.5(3) +1.5 = the t=3 answer right?

OpenStudy (amorfide):

http://prntscr.com/6fy2he or you can use that formula

OpenStudy (anonymous):

and then 1.1(5)^2 -2.5(5) + 1.5 = the t=5 answer?

OpenStudy (amorfide):

\[\frac{ f(3+2)-f(3) }{ 2 }\]

OpenStudy (anonymous):

you mean, 3 +5?

OpenStudy (anonymous):

oh wait nvr mind =) So i have my two answers: 3.9 and 16.5 =)

OpenStudy (amorfide):

no 3 represents the starting x value and the triangle x means the change in x from 3 to 5 so 2=triangle x

OpenStudy (anonymous):

okay that's fine, u r starting to confuse me =) lol

OpenStudy (amorfide):

sorry, I should have mentioned it in terms of t lol

OpenStudy (anonymous):

lol okay =)

OpenStudy (anonymous):

So. . . . . . .? is it 12.6?

OpenStudy (amorfide):

the original expression for a derivative of a function is \[f'(t)=\frac{ f(t _{0} + (t _{1}-t _{0})) - f(t _{0}) }{ t _{1}-t _{0} }\] t0 is the starting t value t=3 t1 is ending t value t=5 rate of change means you want to see how the gradient function also known as the differential function changes

OpenStudy (amorfide):

there you go got there in the end

OpenStudy (amorfide):

so the answer is

OpenStudy (amorfide):

12.6/2

OpenStudy (amorfide):

6.3

OpenStudy (anonymous):

okay awesome thx =)

OpenStudy (amorfide):

sorry for taking so long

OpenStudy (anonymous):

lol it's okay =) thx anyway